2D Kinematics problem

Initial Speed = 3.92 m/s

50 degree Angle

1.045m above the ground

Calculate the expected distance traveled by the ball.
Answer #1

Vox = V.CO= (3.92 m/s) COS 58 Vox = 2.52 m/s Voy = Vosino = 300/5 From kinematic equation. y=yot voyt t ayt 0 = 1.045m +(3197